OpenStudy (anonymous):

The rectangular-shaped landing area of (x3 + 4x2 + 3x + 12) ft2 is shaded. The dimensions of the landing area can be found by factoring. What are the length and width of the landing area? (Assume x > 3.) ft (smaller value) ft (larger value) . . . . .

OpenStudy (anonymous):

First factor: \[ \begin{split} x3 + 4x2 + 3x + 12 &=x^2(x+4)+3x+12\\ &=x^2(x+4) + 3(x+4)\\ &=(x^2+3)(x+4) \end{split} \]